OpenStudy (anonymous):

Trig function help? (Sin, Cos, Tan)? 3. A forest ranger spots a fire from a 28-foot tower. The angle of depression from the tower to the fire is 11 . To the nearest foot, how far is the fire from the base of the tower? Show the steps you use to find the solution.

OpenStudy (ranga):



OpenStudy (anonymous):

So what do I use? @ranga

OpenStudy (ranga):

tan(ABC) = opposite / adjacent = AC / AB tan(11) = 28 / AB AB = 28 / tan(11) = ? (set you calculator to degree and not radians)

OpenStudy (anonymous):

I am using a google calculator xD

OpenStudy (ranga):

then set your Google calculator to degrees.

OpenStudy (ranga):

What are you getting?

OpenStudy (anonymous):

-0.1239

OpenStudy (anonymous):

no wait I got 144.0475

OpenStudy (ranga):

Yes. Round it to the nearest feet. So 144 feet.